Circular No. 9063

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2009 P2 (BOATTINI) A. Boattini reports his discovery of a strongly condensed comet with a

8′′ coma and no tail on four co-added 30-s CCD exposures taken with the Catalina 0.68-m Schmidt telescope (discovery observation tabulated below). Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, other CCD astrometrists have also commented on the objects slight cometary appearance. P. Birtwhistle (Great Shefford, Berkshire, England, 0.40-m reflector, Aug. 15.9 UT) finds the object to be distinctly diffuse with diameter 9′′, possibly elongated east-west (though with no obvious tail); on Aug. 18.0, he found it to be slightly softer than the nearby stellar images (maximum dimensions of 11′′ × 9′′, elongated east-west, with a star of similar brightness have size ≈ 6′′ × 6′′). S. Foglia reports that fifteen stacked 60-s images obtained by R. Holmes (Charleston, IL, USA, 0.61-m astrograph, Aug. 16.2) show the comet to be slightly elongated (size 6′′ × 4′′) in p.a. 222^{o}. R. Ligustri (Talmassons, Udine, Italy, 0.25-m reflector, Aug. 16.3) reports that fifteen 180-s exposures show the object to have a slightly diffuse coma. F. Fratev (Sofia, Bulgaria, 0.35-m reflector, Aug. 18.0) writes that thirty stacked 40-s unfiltered exposures show a coma nearly 15′′ in diameter. E. Guido (Castellammare di Stabia, Italy) reports that G. Sostero and he co-added twenty 120-sunfiltered exposures obtained remotely with a 0.25-m reflector near Mayhill, NM, USA, on Aug. 18.3, revealing an extremely compact coma ∼ 8′′ in diameter, slightly elongated toward the southwest.

2009 UT α_{2000} δ_{2000} Mag. Observer
Aug. 15.44097 23^{h}56^{m}33.^{s}82 +17^{o}28′07′′.8 19.2 Boattini

The available astrometry, the following preliminary parabolic orbital elements by B. G. Marsden, and an ephemeris appear on MPEC 2009-Q14.

T = 2010 July 24.182 TT  ω = 92.^{o}679
Ω = 61.741 }2000.0
q = 6.13366 AU  i = 163.699

COMET C/2006 W3 (CHRISTENSEN) Visual total-magnitude estimates by J. J. Gonzalez, Leon, Spain (0.20-

m reflector unless otherwise noted): Mar. 3.23 UT, 9.6; Apr. 20.17, 9.2; May 2.17, 8.9; June 1.08, 9.0; July 3.08, 8.6; 19.08, 8.6; 25.00, 8.3 (10×50 binoculars); Aug. 11.90, 8.0 (10×50 binoculars).

2009 August 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green